# Exact Model Reduction for Continuous-Time Open Quantum Dynamics

### v2

@@ VI.1 Dissipative central spin model
Even in their most general form involving non-collective and anisotropic couplings, central-spin model Hamiltonians remain formally exactly solvable in the sense of admitting an exact algebraic diagonalization via Bethe Ansatz techniques, as long as appropriate equations are obeyed [99, 100] and factorized initial conditions are assumed. In practice, however, extracting the desired output dynamics from the Bethe Ansatz solution remains very hard, and various (e.g., mean-field) approximations and (or) numerical techniques must be employed [116, 117]. Although the disorder-free case of uniform couplings we consider is highly idealized, it provides a useful playground for showing how observable-based MR may be carried out by using the general approach of Sec. V.3. Having achieved that, in Sec. VI.1.4 we further demonstrate how additional MR reduction can arise, in the sense of Sec. III, depending on the choice of the (reduced) initial condition.
VI.1.1 Non-dissipative central spin with non-dynamical bath
@@
which implies a quadratic scaling [^4] As one may show, a linear scaling of both $\dim(\mathscr{N}^{\perp})$ and $\dim\mathscr{O}$ is found for the class of analytically solvable central-spin models with single-axis coupling mentioned before; see Appendix B.1.. As we noted, projecting onto $\mathscr{N}^{\perp}$ may be fully appropriate for the purposes of simulating the dynamics on a classical computer and numerically computing the output quantities of interest.
@@
In order to explicitly carry out the observable-based quantum MR procedure presented in Sec. V.3, the starting point is to define $\mathcal{H}_{\check{B},j}\equiv{\rm span}\{\left|j,m\right>,\,m=-j,\dots,j\}$, so that $\mathcal{H}_{S}\otimes\mathcal{H}_{\check{B},j}=\mathcal{H}_{F,j}$. We then proceed to derive the reduced state, Hamiltonian, and observables accordingly. While the details are included in Appendix B.2, the final result may be summarized as follows. In the virtual bipartition above, let $W_{j}^{\dagger}$ be the (non-square) isometry that maps the spin-$j$ invariant subspace $\mathcal{H}_{j}=\mathcal{H}_{F,j}\otimes\mathcal{H}_{G,j}$ onto $\mathcal{H}$. Then, the reduced state on $\mathcal{D}(\check{\mathcal{H}})$ can be evaluated as
@@
Similarly, the observables of interest are reduced according to
@@
with the reduced state on the central spin being given by
@@
This property can be useful both to gain qualitative insight into time evolution and to simplify the simulation of the model, as it is possible to analyze and simulate each block in parallel, and then sum the different contributions to obtain the full trajectory; see also Remark 6 in Sec.  VI.1.3.
@@
VI.1.2 Effect of unitary bath dynamics
@@
Figure 3: Comparison between full vs. reduced dynamics for a central-spin model that couples via a general XYZ collective Hamiltonian, Eq. (39), to a bath with Ising intra-bath interactions, Eqs. (49) and (51), for $N=7$. Left: Time-dependent central spin polarization, $\langle\sigma_{z}^{(1)}(t)\rangle$. Right: Time-dependent central spin entropy, $\mathscr{H}[\rho_{S}(t)]$. In both cases, with $\hbar=k_{B}=1$, time is in units of inverse temperature, and the parameters are: $\omega_{1}=0.83$, $\eta=0.01$, $\overline{\omega}=1$, $\mu=0.02$, $A_{x}=4$, $A_{y}=0.1$, $A_{z}=-0.5$, with three different bath-interaction strengths $\lambda$ being used in both the full (dots) and the reduced (solid and dashed lines) models. The initial joint state is $\rho_{0}=\left|0\right>\left<0\right|\otimes\rho_{B}$, with $\rho_{B}$ being the thermal state at inverse temperature $\beta=50$.
@@
However, since the permutation group is non-Abelian, we have $[\mathscr{G}_{N},H^{\text{Heis}}_{B}]\neq 0$, except in the uniform-coupling limit where $B_{ik}\equiv B$. Therefore, $\mathscr{G}_{N}$ remains an ODS despite not being a symmetry, and we can use its commutant algebra to carry out MR. Since $\check{H}_{B,\text{int}}=0$ for ${H}_{B,\text{int}}\in\mathbb{C}\mathscr{G}_{N}$, the resulting reduced model is then still given by Eqs. (46)-(47), upon redefining $\check{H}_{j}$: if we now denote with $\check{H}_{0,j}$ the Hamiltonian in Eq. (VI.1.1), we have
@@
VI.1.3 Effect of Markovian dissipation on bath spins
@@
where $\mathcal{L}_{j}^{\text{coll}}$ can be interpreted as the reduced Lindbladian in the $j$-th subspace, and $\check{H}_{j}$ has the expression in Eq. (VI.1.1). A graphical representation of the Lindbladian operators’ structure for both the original and reduced model is given in Fig. 5.
Figure 6: Left: Time-dependent central spin polarization, $\langle{\sigma_{z}^{(1)}(t)}\rangle$, under collective bath dissipation of variable strength $\Lambda$, in both the full (dots) and reduced (solid curve) models, for $N=6$. As in [104], we choose an XXZ system-bath coupling Hamiltonian, with $A_{x}=A_{y}=4,A_{z}=2$, $\omega_{1}=4.2$, $\eta=0$, $H_{B}=0$, while the bath is at zero temperature, hence $\rho_{0}=\left|1\middle>\middle<1\right|\otimes\left|00\dots 0\middle>\middle<00\dots 0\right|$. The vertical dashed lines indicate the time-to-SS, $\tau_{\text{ss}}$, for the considered values of $\Lambda$, with the shaded blue region showing a $5\%$ deviation from the asymptotic value. Right: Time-to-SS vs. dissipation strength. The dots represent the values obtained from the numerical solution of the reduced model, while the solid curve represents the approximated estimate obtained in Sec. VI.1.4.
@@
Because $\sigma_{+}^{(k)}$ breaks permutation symmetry, $\mathscr{G}_{N}$ is no longer a strong symmetry; however, the form of the dissipator makes it clear that it remains a weak symmetry of the dynamics. This implies that the output algebra onto which we reduce the model remains the same, that is, $\mathscr{O}=\mathbb{C}\mathscr{G}^{\prime}_{N}$. While the Hamiltonian retains the block-diagonal structure of Eq. (VI.1.1), the reduced noise operators, say, $\{\check{L}_{B,h}^{\text{eff}}\}$ are no longer block-diagonal in the basis defined by $U$, hence Eq. (34) (right) does not hold in this case. The resulting structure is illustrated in Fig. 5, top row. Albeit no closed formula is available to compute the reduced operators when off-diagonal terms are present, the reduced noise operators can still be computed numerically by first computing the super-operators $\mathcal{R}$ and $\mathcal{J}$ as in Eq. (21) and (22), and then by determining effective noise operators $\check{L}^{\text{eff}}_{B,h}\in\mathfrak{B}(\check{\mathcal{H}})$, such that
@@
VI.1.4 Reachable analysis: W-state initialization
In the analysis carried out so far, we have only exploited the possibility to reduce the model of interest based on specified observables. In this section, we exemplify the potential of the reachable-based MR presented in Sec. III.
For the sake of illustrating the procedure through an analytical example, let us first assume that the dynamics exhibit a strong symmetry: specifically, we consider a central-spin model subject to collective dissipation as in Sec. VI.1.3, with
$H_{S}=\tfrac{1}{2}\omega_{1}\sigma_{z}^{(1)},\quad H_{B}=0,\quad H_{\text{int}}=H_{\text{XXZ}},\quad L_{B}^{\text{coll}}=\Lambda J_{+}.$
@@ VI.1.4 Reachable analysis: W-state initialization
with $\dim(\mathscr{R})=5$. One can also verify analytically that $\mathscr{R}$ is an algebra of dimension $5$ (irrespective of the bath’s size), has only partial support in the block $\mathfrak{B}(\mathcal{H}_{F,j^{*}})$, and has a Wedderburn decomposition equal to alg$(\mathscr{R})\simeq\mathbb{C}\oplus\mathbb{C}^{2\times 2}\oplus\mathbb{0}_{R}$, where $\dim(\mathcal{H}_{R})=2N-3$. Following the discussion given in Sec. III.2.1, we can restrict our attention to the support of the algebra, so that the reduced model obtained after observable-based and reachable-based MR is going to be defined over $\mathbb{C}\oplus\mathbb{C}^{2\times 2}\subset\mathbb{C}^{3\times 3}$. Since, in addition, we have that $\mathscr{R}={\rm alg}(\mathscr{R})$, we can take $\mathscr{D}=\mathscr{R}$ in the procedure described in Sec. III.3 and proceed to compute the maps $\mathcal{R}$ and $\mathcal{J}$ over the support of $\mathscr{D}$ to obtain the reduced model.
@@
$\tau(t)\equiv\sum_{k,l=0,1}\left|k\middle>\middle<l\right|{\rm tr}\left[\left|k\middle>\middle<l\right|\otimes\left|j^{*}-f(k)\middle>\middle<j^{*}-f(l)\right|\check{\rho}(t)\right]$
@@
VI.2 Boundary-driven XXZ model
@@
We consider a finite, open-boundary chain of $N$ spin-$1/2$ arranged on a line, interacting with each other via a nearest-neighbor XXZ coupling Hamiltonian, of the form
@@ VI.2.1 Symmetries of the model
Figure 8: Dimensions of the observable space $\mathscr{N}^{\perp}$ and output algebra $\mathscr{O}$ compared to the dimensions of the full operator space $\mathfrak{B}(\mathcal{H})$ for the boundary-driven XXZ model. While the dimension of $\mathbb{C}\mathscr{G}_{\varphi}^{\prime}$ is given by the analytical expression in Eq. (59), the dimensions of $\mathscr{N}^{\perp}$ and $\mathscr{O}$ have been computed numerically using the procedure described in Sec. IV.2. The model parameters are $\Gamma=1$, $\Delta=0.5$, $\kappa=1.2$, and $\mu=0.1$.
@@
$\dim(\mathbb{C}\mathscr{G}_{\varphi}^{\prime})=\sum_{k=0}^{N}\binom{N}{k}^{\!2}=\binom{2N}{N}\approx\frac{4^{N}}{\sqrt{N}},$ (59)
@@
$\max_{k}\dim(\mathfrak{B}(\mathcal{H}_{F,k}))=\binom{N}{\lfloor{{N}/{2}}\rfloor}^{2}\approx 2\,\frac{4^{N}}{N},$
@@
To verify whether one may achieve a larger reduction than the one provided by the weak symmetry group $\mathbb{C}\mathscr{G}_{\varphi}^{\prime}$, the observable algebra $\mathscr{O}={\rm alg}(\mathscr{N}^{\perp})$ was numerically determined by implementing the algorithm described in Sec. IV.2. Numerical calculations of the observable space $\mathscr{N}^{\perp}$, the output algebra $\mathscr{O}$, and its Wedderburn decomposition were carried out for chains with up to $N=5$ spins; the results are summarized in Table 1 and Fig. 8. As one may observe from Fig. 8, the dimension $\mathscr{O}$ coincides with the one of $\mathbb{C}\mathscr{G}_{\varphi}^{\prime}$ for up to $N=5$. Moreover, from Table 1, one can observe that the Wedderburn decomposition of $\mathscr{O}$ has precisely $N+1$ blocks of the same dimension as the eigenspaces of $M$. Further numerical testing confirmed that, up to $N=5$ qubits, $\mathscr{O}=\mathbb{C}\mathscr{G}_{\varphi}^{\prime}$. We conjecture this to be true for arbitrary $N$.
@@
VI.2.2 Numerical model reduction and simulations
Having obtained the relevant output algebra $\mathscr{O}$ , let us denote with $U$ the unitary change of basis that decomposes it into its Wedderburn decomposition. In particular, $U$ can be taken as the permutation matrix that orders the eigenvalues of $M$ and groups them together. Thanks to their commutation with $\mathscr{G}_{\varphi}$, both the system’s Hamiltonian and the observables of interest assume a block-diagonal structure in this new basis, with the same block dimensions of the Wedderburn decomposition of $\mathscr{O}$. The four noise operators, instead, assume an upper or lower block-diagonal structure in this basis, as depicted in Fig. 9. As in the case of the central spin model under local dissipation we discussed in Sec. VI.1.3, the off-diagonal terms connect different blocks of $\mathscr{O}$, effectively creating a probability pumping from one block to the other.

### v3

@@ 6.1 Dissipative central spin model
Even in their most general form involving non-collective and anisotropic couplings, central-spin model Hamiltonians remain formally exactly solvable in the sense of admitting an exact algebraic diagonalization via Bethe Ansatz techniques, as long as appropriate equations are obeyed [101, 102] and factorized initial conditions are assumed. In practice, however, extracting the desired output dynamics from the Bethe Ansatz solution remains very hard, and various (e.g., mean-field) approximations and (or) numerical techniques must be employed [118, 119]. Although the disorder-free case of uniform couplings we consider is highly idealized, it provides a useful playground for showing how observable-based MR may be carried out by using the general approach of Sec. 5.3. Having achieved that, in Sec. 6.1.4 we further demonstrate how additional MR reduction can arise, in the sense of Sec. 3, depending on the choice of the (reduced) initial condition.
6.1.1 Non-dissipative central spin with non-dynamical bath
@@
which implies a quadratic scaling [^4] As one may show, a linear scaling of both $\dim(\mathscr{N}^{\perp})$ and $\dim\mathscr{O}$ is found for the class of analytically solvable central-spin models with single-axis coupling; see Appendix B.1.. As we noted, projecting onto $\mathscr{N}^{\perp}$ may be fully appropriate for the purposes of simulating the dynamics on a classical computer and numerically computing the output quantities of interest.
@@
In order to explicitly carry out the observable-based quantum MR procedure presented in Sec. 5.3, the starting point is to define $\mathcal{H}_{\check{B},j}\equiv{\rm span}\{\left|j,m\right>$, with $m=-j,\dots,j\}$, so that $\mathcal{H}_{S}\otimes\mathcal{H}_{\check{B},j}=\mathcal{H}_{F,j}$. We then proceed to derive the reduced state, Hamiltonian, and observables accordingly. While the details are included in Appendix B.2, the final result may be summarized as follows. In the virtual bipartition above, let $W_{j}^{\dagger}$ be the (non-square) isometry that maps the spin-$j$ invariant subspace $\mathcal{H}_{j}=\mathcal{H}_{F,j}\otimes\mathcal{H}_{G,j}$ onto $\mathcal{H}$. Then, the reduced state on $\mathcal{D}(\check{\mathcal{H}})$ can be evaluated as
@@
Similarly, the observables of interest are reduced according to the following:
@@
with the reduced state on the central spin given by
@@
This property can be useful both to gain qualitative insight into time evolution and to simplify the simulation of the model, as it is possible to analyze and simulate each block in parallel, and then sum the different contributions to obtain the full trajectory; see also Remark 6 in Sec.  6.1.3.
@@
6.1.2 Effect of unitary bath dynamics
@@
Figure 3: Comparison between full vs. reduced dynamics for a central-spin model that couples via a general XYZ collective Hamiltonian, Eq. (39), to a bath with Ising intra-bath interactions, Eqs. (49) and (51), for $N=7$. Left: Time-dependent central-spin polarization, $\langle\sigma_{z}^{(1)}(t)\rangle$. Right: Time-dependent central-spin entropy, $\mathscr{H}[\rho_{S}(t)]$. In both cases, with $\hbar=k_{B}=1$, time is in units of inverse temperature, and the parameters are: $\omega_{1}=0.83$, $\eta=0.01$, $\overline{\omega}=1$, $\mu=0.02$, $A_{x}=4$, $A_{y}=0.1$, $A_{z}=-0.5$. Three different bath-interaction strengths $\lambda$ are used in both the full (dots) and the reduced (solid and dashed lines) models. The initial joint state is $\rho_{0}=\left|0\right>\left<0\right|\otimes\rho_{B}$, with $\rho_{B}$ being the thermal state at inverse temperature $\beta=50$.
@@
However, since the permutation group is non-Abelian, we have $[\mathscr{G}_{N},H^{\text{Heis}}_{B}]\neq 0$, except in the uniform-coupling limit where $B_{ik}\equiv B$. Therefore, $\mathscr{G}_{N}$ remains an ODS despite not being a symmetry, and we can use its commutant algebra to carry out MR. Since $\check{H}_{B,\text{int}}=0$ for ${H}_{B,\text{int}}\in\mathbb{C}\mathscr{G}_{N}$, the resulting reduced model is then still given by Eqs. (46)-(47), upon redefining $\check{H}_{j}$: if we now denote with $\check{H}_{0,j}$ the Hamiltonian in Eq. (6.1.1), we have
@@
6.1.3 Effect of Markovian dissipation on bath spins
@@
where $\mathcal{L}_{j}^{\text{coll}}$ can be interpreted as the reduced Lindbladian in the $j$-th subspace, and $\check{H}_{j}$ has the expression in Eq. (6.1.1). A graphical representation of the Lindbladian operators’ structure for both the original and reduced model is given in Fig. 5.
Figure 6: Left: Time-dependent central spin polarization, $\langle{\sigma_{z}^{(1)}(t)}\rangle$, under collective bath dissipation of variable strength $\Lambda$, in both the full (dots) and reduced (solid curve) models, for $N=6$. As in [106], we choose an XXZ system-bath coupling Hamiltonian, with $A_{x}=A_{y}=4,A_{z}=2$, $\omega_{1}=4.2$, $\eta=0$, $H_{B}=0$, while the bath is at zero temperature, hence $\rho_{0}=\left|1\middle>\middle<1\right|\otimes\left|00\dots 0\middle>\middle<00\dots 0\right|$. The vertical dashed lines indicate the time-to-SS, $\tau_{\text{ss}}$, for the considered values of $\Lambda$, with the shaded blue region showing a $5\%$ deviation from the asymptotic value. Right: Time-to-SS vs. dissipation strength. The dots represent the values obtained from the numerical solution of the reduced model, while the solid curve represents the approximated estimate obtained in Sec. 6.1.4.
@@
Because $\sigma_{+}^{(k)}$ breaks permutation symmetry, $\mathscr{G}_{N}$ is no longer a strong symmetry; however, the form of the dissipator makes it clear that it remains a weak symmetry of the dynamics. This implies that the output algebra onto which we reduce the model remains the same, that is, $\mathscr{O}=\mathbb{C}\mathscr{G}^{\prime}_{N}$. While the Hamiltonian retains the block-diagonal structure of Eq. (6.1.1), the reduced noise operators, say, $\{\check{L}_{B,h}^{\text{eff}}\}$ are no longer block-diagonal in the basis defined by $U$, hence Eq. (34) (right) does not hold in this case. The resulting structure is illustrated in Fig. 5, top row. Albeit no closed formula is available to compute the reduced operators when off-diagonal terms are present, the reduced noise operators can still be computed numerically by first computing the super-operators $\mathcal{R}$ and $\mathcal{J}$ as in Eq. (21) and (22), and then by determining effective noise operators $\check{L}^{\text{eff}}_{B,h}\in\mathfrak{B}(\check{\mathcal{H}})$, such that
@@
6.1.4 Reachable analysis: W-state initialization
In the analysis carried out so far, we have only exploited the possibility to reduce the model of interest based on specified observables. In this section, we exemplify the potential of the reachable-based MR presented in Sec. 3.
For the sake of illustrating the procedure through an analytical example, let us first assume that the dynamics exhibit a strong symmetry: specifically, we consider a central-spin model subject to collective dissipation as in Sec. 6.1.3, with
$H_{S}=\tfrac{1}{2}\omega_{1}\sigma_{z}^{(1)},\,\,\,H_{B}=0,\,\,\,H_{\text{int}}=H_{\text{XXZ}},\,\,\,L_{B}^{\text{coll}}=\Lambda J_{+}.$
@@ 6.1.4 Reachable analysis: W-state initialization
with $\dim(\mathscr{R})=5$. One can also verify analytically that $\mathscr{R}$ is an algebra of dimension $5$ (irrespective of the bath’s size), has only partial support in the block $\mathfrak{B}(\mathcal{H}_{F,j^{*}})$, and has a Wedderburn decomposition equal to alg$(\mathscr{R})\simeq\mathbb{C}\oplus\mathbb{C}^{2\times 2}\oplus\mathbb{0}_{R}$, where $\dim(\mathcal{H}_{R})=2N-3$. Following the discussion given in Sec. 3.2.1, we can restrict our attention to the support of the algebra, so that the reduced model obtained after observable-based and reachable-based MR is going to be defined over $\mathbb{C}\oplus\mathbb{C}^{2\times 2}\subset\mathbb{C}^{3\times 3}$. Since, in addition, we have that $\mathscr{R}={\rm alg}(\mathscr{R})$, we can take $\mathscr{D}=\mathscr{R}$ in the procedure described in Sec. 3.3 and proceed to compute the maps $\mathcal{R}$ and $\mathcal{J}$ over the support of $\mathscr{D}$ to obtain the reduced model.
@@
$\tau(t)\equiv\sum_{k,l=0}^{1}\left|k\middle>\middle<l\right|{\rm tr}\left[\left|k\middle>\middle<l\right|\otimes\left|j^{*}-f(k)\middle>\middle<j^{*}-f(l)\right|\check{\rho}(t)\right]$
@@
6.2 Boundary-driven XXZ model
@@
We consider a finite, open-boundary chain of $N$ spin-$1/2$ on a line, interacting with each other via a nearest-neighbor XXZ coupling Hamiltonian, of the form
@@ 6.2.1 Symmetries of the model
Figure 8: Dimensions of the observable space $\mathscr{N}^{\perp}$ and output algebra $\mathscr{O}$ compared to the dimensions of the full operator space $\mathfrak{B}(\mathcal{H})$ for the boundary-driven XXZ model.
While the dimension of $\mathbb{C}\mathscr{G}_{\varphi}^{\prime}$ is given by the analytical expression in Eq. (59), the dimensions of $\mathscr{N}^{\perp}$ and $\mathscr{O}$ have been computed numerically using the procedure described in Sec. 4.2. The model parameters are $\Gamma=1$, $\Delta=0.5$, $\kappa=1.2$, and $\mu=0.1$.
@@
$\dim(\mathbb{C}\mathscr{G}_{\varphi}^{\prime})=\sum_{k=0}^{N}\binom{N}{k}^{\!2}=\binom{2N}{N}\approx\frac{4^{N}}{\sqrt{\pi N}},$ (59)
@@
$\max_{k}\dim(\mathfrak{B}(\mathcal{H}_{F,k}))=\binom{N}{\lfloor{{N}/{2}}\rfloor}^{2}\approx 2\,\frac{4^{N}}{\pi N},$
@@
To verify whether one may achieve a larger reduction than the one provided by the weak symmetry group $\mathbb{C}\mathscr{G}_{\varphi}^{\prime}$, the observable algebra $\mathscr{O}={\rm alg}(\mathscr{N}^{\perp})$ was numerically determined by implementing the algorithm described in Sec. 4.2. Numerical calculations of the observable space $\mathscr{N}^{\perp}$, the output algebra $\mathscr{O}$, and its Wedderburn decomposition were carried out for chains with up to $N=5$ spins; the results are summarized in Table 1 and Fig. 8. As one may observe from Fig. 8, the dimension $\mathscr{O}$ coincides with the one of $\mathbb{C}\mathscr{G}_{\varphi}^{\prime}$ for up to $N=5$. Moreover, from Table 1, one can observe that the Wedderburn decomposition of $\mathscr{O}$ has precisely $N+1$ blocks of the same dimension as the eigenspaces of $M$. Further numerical testing confirmed that, up to $N=5$ qubits, $\mathscr{O}=\mathbb{C}\mathscr{G}_{\varphi}^{\prime}$. We conjecture this to be true for arbitrary $N$.
@@
6.2.2 Numerical model reduction and simulations
Having obtained the relevant output algebra $\mathscr{O}$ , let us denote with $U$ the unitary change of basis that decomposes it into its Wedderburn decomposition. In particular, $U$ can be taken as the permutation matrix that orders the eigenvalues of $M$ and groups them together. Thanks to their commutation with $\mathscr{G}_{\varphi}$, both the system’s Hamiltonian and the observables of interest assume a block-diagonal structure in this new basis, with the same block dimensions of the Wedderburn decomposition of $\mathscr{O}$. The four noise operators, instead, assume an upper or lower block-diagonal structure in this basis, as depicted in Fig. 9. As in the case of the central spin model under local dissipation we discussed in Sec. 6.1.3, the off-diagonal terms connect different blocks of $\mathscr{O}$, effectively creating a probability pumping from one block to the other.